OpenStudy (anonymous):

I need help with these algebra 2 questions. 1. What is the vertex form of the equation? y=-x^2+12x-4 2. What is the expression in factored form? 2x^2+16x+24

OpenStudy (anonymous):

Factored Form To do this all you need to do is this First reorder the terms to look like this :) \(24 ~+ ~16x ~+~ 2x^2\) Next you factor out the greatest common factor of 24, 16, and 2, and that is 2, so divide them all by 2 \(24/2 ~+ ~16x/2 ~+~ 2x^2/2\) \(2(12 ~+ ~8 ~+~ x^2)\) Now we have \(2(12 ~+ ~8x ~+~ x^2)\), so next we factor a trinomial (a polynomial with three terms) \(2((6 ~+ ~x)(2 ~+~ x))\) And take away the extra parenthesis and you have \(2(x~+~6)(x~+~2)\)

OpenStudy (anonymous):

Sorry I can't help with the first one

OpenStudy (shinalcantara):

\[y = -x^2 + 12x - 4\] -------------------- Vertex form is given by the equation \[y = a(x-h)^2 + k\] -------------------- \[y = -x^2 + 12x - 4\] group through parenthesis \[y = (-x^2 + 12x) - 4\] factor out -1 \[y = -1(x^2 - 12x) - 4\] Complete the square by adding and subtracting 36 on the right side of the equation \[y = -1(x^2 - 12x+36) - 4 + 36\] Notice that what you'll see are all positive 36 since when you try to distribute -1 outside the quantity it would give you -36 \[y = -1(x-6)^2 +32\]

OpenStudy (anonymous):

Thank you both for helping, I really appreciate it. :)